[Brahms-l] Shift report 20040130 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Fri Jan 30 2004 - 08:07:55 EST
------------------------------------------------------------
Date: Fri Jan 30 08:07:55 2004, Shift: 00:00-08:00
Supervisor: EJ, with: 
Log book#15 pages 45 - 45

Shift Summary:
------------------------------------------------------------
1. We had some good moments with stable beam.  

2. I finished up the runs for FS at 8deg B1219 field.
I got 11K events for the first zero field run.

3. I had one detector trip, TPM1.  I was NOT collecting data at the time, and I got it reset with out much trouble.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 30 08:08:25 2004

This archive was generated by hypermail 2.1.8 : Fri Jan 30 2004 - 08:08:41 EST